# General relativity and the double-zero eigenvalue

### v1

@@ 5.4 Evolution of event horizons
The problem of studying crease structures defined as endpoints of at least two horizon generators on the event horizon of a generic black hole as 2-dimensional spacelike submanifolds is here reformulated as an evolution problem of the crease submanifolds governed by a new dynamical system, the crease flow [21]. This replaces the usual Hamiltonian flow wherein the crease sets appear as its steady-state (i.e., fixed point) solutions, while their bifurcations describe the possible types and topological transformations of the singularities very accurately.
The crease flow introduced in [21] deals with the dynamics of event horizons at caustic points where the usual Taylor expansions used in the literature break down, and consequently it becomes unclear how to make physical sense and describe the evolution that develops such caustics. This has a number of benefits over previous approaches to the problem.
We may write down the defining equations of a wave front for this problem (obtained from the event horizon by extending the generators beyond their past end points as far as possible) as the union of two null hypersurfaces in spacetime intersecting transversally.
The intersections are given by the equations $u-b_{AB}x^{A}x^{B}=0,\,v-c_{AB}x^{A}x^{B}=0$, where $(u,v)$ are null coordinates, $A,B=1,2$, $x,y$, are normal coordinates (the ‘transverse’ phase space), and we expand in normal coordinates up to second order terms with the $b,c$ being quadratic polynomials in $x,y$.
We can use suitably defined 2-spheres of radius $\lambda=\sqrt{uv-T^{2}}$, $T$ being a time function (synchronous coordinate) in the place of the null coordinates so that to any given point $(u,v)$ there corresponds a value of $\lambda$, and the $\lambda$-lines screen the $(u,v)$-plane. We then introduce the crease flow for the evolution of the crease sets as that governed by the dynamical system,
$\begin{split}\dot{x}&=\lambda+b_{AB}x^{A}x^{B}:\quad g_{1},\\ \dot{y}&=\lambda+c_{AB}x^{A}x^{B}:\quad g_{2},\end{split}$ (5.1)
@@
The correct nondegeneracy and transversality conditions on the crease flow imply that the crease evolution is a bifurcation problem of codimension three. This result requires the deployment of several notions of singularity theory into this problem (for instance the nature of the Jacobian of the crease flow at the origin).
Further, the bifurcation sequence developed in Section 4 can be applied to the crease flow leading to three different normal forms and corresponding versal unfoldings (each of which having one distinguished parameter, the $\lambda$, and three unfolding parameters). These results depend on the nature of the zero set and the sign of the discriminant of the Jacobian, making the present problem perhaps the most complicated among all those analysed so far (with the possible exception of the Friedmann-Lemaître problem - see below).
We may depict the resulting situation in projections where two of the unfolding parameters are set to zero. In this representation, the problem is described by the bifurcation diagram shown in Fig. 5 we met earlier in shifted variables $X,Y$ and parameters $\mu_{1},\mu_{2}$ (one of them is the $\lambda$ and the other one of the unfolding parameters, cf. [21], Eq. (14)).
In the bifurcation diagram of Fig. 5, we see the resulting ‘liquid-like’ picture of evolving, in principle observable, bifurcating caustics on the event horizon corresponding to intersections of null hypersurfaces in spacetime. This is based on the joined effects of the three possible bifurcations in this problem, namely, saddle-node when crossing any of the two axes, pitchfork when crossing the parabola, and Hopf when crossing the positive $\mu_{1}$-axis, acting on the following caustics: swallowtail, folded Whitney, and a pair of ‘embedded in each other’ (i.e., combined) Whitney singularities, all three being steady states of the crease flow. Using $z$ as a third coordinate, the latter are given by the forms  $x^{2}=zy^{2}$ (standard Whitney), and $x^{2}=z^{3}y^{2}$ (folded).
As the parameter changes differently at different spacetime points, these effects occur simultaneously, thus offering a very rich picture for the evolution of the crease sets. We also note the formation of the stable isolated limit cycle in the $\alpha$-stratum attracting nearby orbits and disappearing upon crossing the $\nu$-line into the $\eta$-stratum.
@@ 5.5 The Friedmann-Lemaître equations
Figure 6: The axis of the parameter $\gamma$. Away from the two shown values of $\gamma$, the system can be either time-symmetric or time-asymmetric. In either case, a passage through the time-(a)symmetric value $-2/3$ (resp. $2/3$) dualizes the behaviour of a time-(a)symmetric that existed for the system when $\gamma\neq\pm 2/3$.
@@ 5.5.3 The FL cusp
and may also be found in, for instance, [6], pp. 444-5, for the (+)(+)( + )-case, and [9], p. 358 or [46], p. 1038, Fig. 15 for ()(-)( - )-case.
@@ 5.5.4 Codimension-3
can be found in [46], figs. 3, 4 for the saddle case, figs. 7, 8 for the focus case, and figs. 13, 14 for the elliptic case (see also [47]).
This is the dynamically richest case among all versally unfolded FL cosmologies (for more details, see [22]).
@@ 5.5.5 Some remarks
Suppose that the system is found itself in one of the five possible ‘ranges’ of $\gamma$-values as in Fig. 6 (the two point values are also regarded as two of these ranges). Then the consequent behaviour of the system is completely determined by the unfolding parameters $\mu_{i},i=1,2$ (and also $\mu_{3}$ for the codimension-3 unfolding (5.10)), in the sense that it moves from one bifurcation diagrams to the next as $\gamma$ varies, and from phase portrait to the next as the unfolding parameters $\mu$ vary, all these transfigurations being smooth.
@@ 6 Discussion
This work describes the path from the large body of known and well-understood results related to hyperbolic-like approaches in general relativity to a consistent structurally stable framework, a new but yet largely unknown structure that emerges when unfolding general relativistic systems as the five ones described in this paper. This ‘theory’ eventually contains all versal unfoldings of general relativity.
Our approach to this problem is based on fundamental mathematical works in the subject of bifurcation theory, starting by H. Poincaré around the 1890, developed further by A. A. Andronov, R. Thom, V. I. Arnold, later by F. Takens, R. I. Bogdanov, and more recently by F. Dumortier, H. Zoladek, and many other mathematicians.
In this connection, we make the following remarks. Consider the Einstein equations with a cosmological constant $\Lambda$ considered as a parameter in the sense advanced in this paper. This is clearly an unfolding of the vacuum Einstein equations, but is it a versal unfolding of them [^5] An answer to this question could perhaps replace the question mark at the end of the vacuum equations in a famous photo!? According to the work in this paper, the answer to this question appears to be negative. We have shown that the Einstein equation with $\Lambda$ contains symmetric subsystems, e.g., the FL equations (with $\gamma=0$ in this case), for which the versal unfoldings have codimension greater than one, and three at most. Therefore the Einstein-$\Lambda$ equations have to be versally unfolded and their codimension appears to be at least three. We conjecture that the codimension of the Einstein-$\Lambda$ equations (without matter) is exactly equal to three.
@@
This theory cannot correspond to one with additional action terms, because of the following observation. Consider the FRW problem discussed earlier in Section 5. In that problem, we start from the well-known results about the behaviour of the simplest FRW models and end up with their versal unfolding, namely, the equation $\Omega^{\prime}=\Omega(\Omega-1)+\nu$, where $\nu$ is the unfolding parameter. This is a perturbation of the Friedmann equations via adding lower-order terms (lower with respect to the first nonzero term in the Taylor expansion), and it is versal i.e., all terms of $O(3)$ or higher have no effect on the nature of the bifurcations. As we discussed, the solutions coming from the versal unfolding equation have properties not met in the standard theory and emerge only as bifurcations. In this sense they represent not modifications of the standard theory but a certain completion of it at a deeper level (as we already discussed earlier), one that is very difficult to capture without bifurcation theory. This way gravitational bifurcation theory may considerably expand the present horizon of gravity research.
@@
Obviously one may consider unfolding a theory that contains higher-order curvature terms or is of higher-dimensionality than four, for example in the framework of some modified gravity theory of effective string theory. Although this would lead to interesting projects in gravitational bifurcation theory, we note that there are already some answers to such a problem stemming from the results discussed in this paper. For example, suppose that one considers a Friedmann universe with a perfect fluid with or without a $\Lambda$ in a modified gravity or string theoretic context. If the linear part of the resulting equations turns out to be the same as the corresponding ones discussed here, then the answer will be identical to the ones we found here. This will be so due to the power of the versal unfolding construction. However, the inclusion of other matter fields, for instance a scalar field component, will drastically alter the conclusions and possibly increase the dimensionality of the dynamical systems and/or the codimension of the unfoldings (even in cases with identical symmetries as the ones here, for instance $\mathbb{Z}_{2}$).
@@
Amongst the many fundamental issues emerging in the framework of gravitational bifurcation theory and deserve a comment, we shall briefly discuss only two. Firstly, since the two appear intertwined in any bifurcation diagram, what is the relation between the unfolding parameter (‘motion’ in parameter space) and the time (evolution in phase space)? In particular, which of the two $\mu$ or $t$ is prior? Unconstrained motion in parameter space appears to be a cause for the appearance of a ‘time succession’ in phase space. Consider evolution in a stratum without equilibria, e.g., the negative axis in Fig. 1 right, or the $\chi$-stratum in Fig. 4, and the motion of a parameter point around the origin in parameter space. Passing through a bifurcation set changes the number and nature of equilibria and forces the phase point to move on a new orbit in the ‘next’ phase portrait. Such a smooth transition of the phase point between different phase portraits introduces a time progression into the problem associated with a global change (i.e., the appearance of one or more new equilibria). This state of affairs begs the question as to what causes motion in parameter space? We believe that an answer to this question is related to the proven stability of the versal unfolding. In other words, the system collapses when the stability of the whole construction is disrupted, something which as we have shown cannot really happen in a versal unfolding.

### v2

@@ 3.5 Evolution of event horizons
The problem of studying crease structures defined as endpoints of at least two horizon generators on the event horizon of a generic black hole as 2-dimensional spacelike submanifolds may be reformulated as an evolution problem of the crease submanifolds governed by a new dynamical system, the crease flow [23]. This replaces the usual Hamiltonian flow and the crease sets appear as its steady-state (i.e., fixed point) solutions while their bifurcations describe the possible types and topological transformations of the singularities very accurately.
The crease flow introduced in [23] deals with the dynamics of event horizons at caustic points where the usual Taylor expansions break down, and consequently it becomes unclear how to make physical sense and describe the evolution that develops such caustics. This approach has a number of benefits over previous approaches to the problem. We may write down the defining equations of a wave front for this problem (obtained from the event horizon by extending the generators beyond their past end points as far as possible) as the union of two null hypersurfaces in spacetime intersecting transversally. The intersections are given by the equations $u-b_{AB}x^{A}x^{B}=0,\,v-c_{AB}x^{A}x^{B}=0$, where $(u,v)$ are null coordinates, $A,B=1,2$, $x,y$, are normal coordinates (the ‘transverse’ phase space), and we expand in normal coordinates up to second order terms with the $b,c$ being quadratic polynomials in $x,y$.
We can then use suitably defined 2-spheres of radii $\lambda=\sqrt{uv-T^{2}}$, $T$ being a time function (synchronous coordinate) in the place of the null coordinates so that to any given point $(u,v)$ there corresponds a value of $\lambda$, and the $\lambda$-lines screen the $(u,v)$-plane. We then introduce the crease flow for the evolution of the crease sets as that governed by the dynamical system,
$\begin{split}\dot{x}&=\lambda+b_{AB}x^{A}x^{B},\\ \dot{y}&=\lambda+c_{AB}x^{A}x^{B},\end{split}$ (3.11)
@@
The correct nondegeneracy and transversality conditions on the crease flow imply that the crease evolution is a bifurcation problem of codimension three. This result requires the deployment of several notions of singularity theory into this problem (for instance the nature of the Jacobian of the crease flow at the origin). Further, one may show that the long bifurcation sequence developed in Section 2 can be applied to the crease flow problem. This leads in particular to three different normal forms and corresponding versal unfoldings (each of which having one distinguished parameter, the $\lambda$, and three unfolding parameters). The results depend on the nature of the zero set and the sign of the discriminant of the Jacobian, making the present problem perhaps the most complicated among all those analysed so far (with the possible exception of the Friedmann-Lemaître problem - see below).
We may depict the resulting situation in projections where two of the unfolding parameters are set to zero. In this representation, the problem is described by the bifurcation diagram shown in Fig. 5 we met earlier in shifted variables $X,Y$ and parameters $\mu_{1},\mu_{2}$ (one of them is the $\lambda$ and the other one of the unfolding parameters, cf. [23], Eq. (14)). In the bifurcation diagram of Fig. 5, we see the resulting ‘liquid-like’ picture of evolving, in principle observable, bifurcating caustics on the event horizon corresponding to intersections of null hypersurfaces in spacetime. This is based on the joined effects of the three possible bifurcations in this problem, namely, saddle-node when crossing any of the two axes, pitchfork when crossing the parabola, and Hopf when crossing the positive $\mu_{1}$-axis, acting on the following caustics: swallowtail, folded Whitney, and a pair of ‘embedded in each other’ (i.e., combined) Whitney singularities, all three being steady states of the crease flow. Using $z$ as a third coordinate, the latter are given by the forms  $x^{2}=zy^{2}$ (standard Whitney), and $x^{2}=z^{3}y^{2}$ (folded).
As the parameter changes differently at different spacetime points, these effects occur simultaneously, thus offering a very rich picture for the evolution of the crease sets. We also note the formation of the stable isolated limit cycle in the $\alpha$-stratum of Fig. 5 attracting nearby orbits and disappearing upon crossing the $\nu$-line into the $\eta$-stratum.
@@ 3.6 The Friedmann-Lemaître equations
Figure 6: The axis of the parameter $\gamma$ which is responsible for the amazing variety of behaviours stemming from the FL equations as $\gamma$ moves along this axis and jumps from one bifurcation diagram to the next passing gradually from all nine possible such diagrams determined by the $\Lambda$ and the various unfolding parameters necessary in each case. At the two $\gamma$ values ±$\pm 2/3$, the bifurcation diagrams are qualitatively equivalent to those shown previously in Figs. 2-5. Away from the two shown values of $\gamma$, the system can be either time-symmetric (i.e., the ‘FL cubic’) or time-asymmetric (i.e., the ‘FL cusp’). In either case, a passage through the time-(a)symmetric value $-2/3$ (resp. $2/3$) dualizes (i.e., flips) the behaviour of a time-(a)symmetric that existed for the system when $\gamma\neq\pm 2/3$ to the opposite one.
@@ 3.6.3 The FL cusp
cf. Figs. 6, 7 of [24]. (They may also be found in, for instance, [61], pp. 444-5, for the (+)(+)( + )-case, and [64], p. 358 or [91], p. 1038, Fig. 15 for ()(-)( - )-case.) For example, the phase portrait corresponding to the origin of the bifurcation diagram is shown in Fig. 7 (this is the zero-parameter case, cf. Ref. [24]).
Figure 7: The cusp-like phase portrait for the zero parameter limit of the versal family (3.19). While a general feature of the quadratic Bogdanov-Takens bifurcation involved in the present FL cusp case is the sequence: saddle-node to Hopf (cycle creation) to homoclinic (cycle destruction), the shown cusp-like phase portrait corresponds to the simplest stratum at the origin: Two special trajectories (the stable and unstable separatrices) form the pointed ‘V’-shape of the semi-cubical parabola; inside that cusp region trajectories slow down as they approach the origin along one branch, linger near the singularity, then are carried away along the other branch. Outside the cusp, orbits simply sweep past, curving around the outside of the cusp separatrices and escaping to or arriving from infinity. The shown cusp-like features are new and correspond to the normal form of the problem at the origin (qualitatively different from the original FL equation).
@@ 3.6.4 Codimension-3
can be found in [91], figs. 3, 4 for the saddle case, figs. 7, 8 for the focus case, and figs. 13, 14 for the elliptic case (see also [92]). The bifurcation diagram has a rhombic structure with ‘vertical lips’ forming two intersecting cusps as in Fig. 8.
Figure 8: The grey area depicts the rhombic structure of the parameter space partitioning it into an internal and an external region for the codimension-3 problem (3.20). The bifurcation diagrams in all cases are very complex and dependent on specific features that appear in each of the three cases of the problem. But there is a common theme in all three cases, namely, the existence of the cuspidal ‘saddle-node’ curves shown present in the rhombic region: the condition for the equilibria of the system (3.20) is the cubic equation, $\mu_{1}+\mu_{2}x\pm x^{3}=0$. For the (+)(+)( + )-case, the discriminant is $\Delta_{+}=-4\mu_{2}^{3}-27\mu_{1}^{2}$, while $\Delta_{-}=4\mu_{2}^{3}-27\mu_{1}^{2}$ for the ()(-)( - )-case in the cubic equation. The two bifurcation sets, $\Delta_{\pm}=0$, for the two pairs of cusps, stratify the parameter space into an internal region where there are three real roots to both cubic equations (and so in all three cases: saddle, focus, and elliptic), and an external region otherwise. The gray ‘vertical-lips’ region in between is exactly the set of $(\mu_{1},\mu_{2})$ where both cubics have three real roots (i.e. $\Delta_{+}\leq 0$ and $\Delta_{-}\geq 0$), and form the cusp-shaped, ‘rhombic’, region in parameter space shown in this figure.
This is the dynamically richest case among all versally unfolded FL cosmologies (cf. [24], and the references given earlier for more details).
@@ 3.6.5 Some remarks
Suppose that the system is found itself in one of the five possible ‘ranges’ of $\gamma$-values as in Fig. 6 (the two point values are also regarded as two of these ranges). Then the consequent behaviour of the system is completely determined by the unfolding parameters $\mu_{i},i=1,2$ (and also $\mu_{3}$ for the codimension-3 unfolding (3.20)), in the sense that it moves from one bifurcation diagram to the next as $\gamma$ varies, and from phase portrait to the next as the unfolding parameters $\mu$ vary, all these transfigurations being smooth.
@@
3.7 Some open problems
The following represents a slice of some open research directions and projects to be followed once the reader feels at home with the material discussed in previous sections of this paper. These projects will deploy and/or use the results of the application of the bifurcation sequence of Section 2. This is not meant to be an end in itself but rather a means to solve physical problems and approach conjectures that depend on (or arise from) an initial treatment of some set of governing gravitational equations. Gravitational bifurcation problems are also beneficial for the bifurcation methods themselves because of the essentially nonlinear and fundamentally different nature of the gravitational equations as compared to other areas of theoretical physics.
  1. 1.
    Physical effects for Friedmann-Lemaître cosmologies and their metamorphoses.
    The extension of Friedmann bifurcations to models with a cosmological constant as discussed earlier presents certain subtle issues which have no counterpart in the simplest models. The main change from a bifurcation theory point of view is that the problem becomes of codimension three (rather than one) and that the cosmological constant is the true bifurcation parameter (rather than the fluid parameter). There are various physical consequences of this result, especially those related to the FL cubic of codimension three and the purpose of this project is to substantiate and further elaborate on certain physical implications that have no counterpart in the original approach to the FL equations and their solutions.
  2. 2.
    Cosmic acceleration and Friedmann cosmology.
    The new Friedmann solutions shown in Fig. 1b lead to accelerating solutions that satisfy all energy conditions. This result may be applied to various problems that are usually attacked through dark energy in an effort to see if they could represent a reliable alternative.
  3. 3.
    Metamorphoses and singularities.
    The physics of the versal unfoldings of singularities and black holes as in Fig. 4 may be developed to explore models which rely on evolution through metamorphoses rather than on the barrier of spacetime singularities. A qualitative comparison of these two aspects of evolution of gravitating systems may be explored in a variety of contexts.
  4. 4.
    Relativistic causality and versality.
    Causal structure techniques may be deployed to describe generic aspects of spacetime bifurcations. These include the development of ‘parametric causal structure theory’, for instance, the parameter dependence of the boundary of the future of a trapped surface during the crossing of a bifurcation set, as in the bifurcation evolution $\chi\to\epsilon_{+}\to o$ in Fig. 4.
  5. 5.
    Acausal versal null hypersurfaces.
    The crease flow evolution may be calculated for a variety of event horizons and its acausal character may be checked in all cases. A very detailed analysis of the physics of all versal unfoldings could then be carried out.
  6. 6.
    Massive scalar fields, the measure problem, and the versal extension.
    The massive scalar field model used in the construction of a measure in the set of all universes may now be revisited. This represents mathematically an exemplary problem in the so-called Hopf-steady state bifurcation. The possibility of a finite, parameter-dependent, measure for the Gibbons-Hawking-Stewart construction and related works may then be fully explored.
  7. 7.
    Transfigurations in scalar field cosmologies with exponential potentials.
    The highly nontrivial mathematical problem of a homogeneous and isotropic scalar field with an exponential potential (which has been studied in many papers) may be explored alternatively by applying the bifurcation sequence, in an effort to discuss the effects of degeneracies on a number of related issues such as the inflationary attractor, the existence of singularities, the nature of initial states, the existence and importance of scaling solutions, as well as various extensions to multiple scalar fields.
  8. 8.
    Bifurcations of Bianchi cosmologies of class A.
    The Wainwright-Hsu equations, which form the basis of many studies in Bianchi models, possess a kind of dihedral symmetry which makes the deployment of bifurcation methods very interesting (cf. [86]). The problem of the description of Bianchi models through their metamorphoses has not been addressed to any systematic degree so far. We expect to explore a number of fundamental attributes of anisotropic cosmology from this or similar viewpoints which will hopefully shed new light to a number of basic questions.
  9. 9.
    Hamiltonian cosmological bifurcations.
    There are many problems in cosmology which admit a Hamiltonian formulation, especially anisotropic models. The theory of Hamiltonian normal forms is of course a time-honoured subject, but to the best of our knowledge has not been applied in this framework before, and the same is true for Hamiltonian bifurcations. There are a number of interesting problems in the framework which may be studied in an effort to address some of the many conjectures in this field.
  10. 10.
    Cosmological perturbations and bifurcation theory.
    Using covariant methods in conjunction with, for instance, the Wainwright-Hsu equations, one is able to write down autonomous systems for many cosmological density perturbations problems. Although the stability of some cases has been studied, the methods are solely based on hyperbolic analysis. A program to revisit this whole area and apply the bifurcation sequence in an effort to decide which of the conclusions of this field remain true in this more general case may therefore be worthwhile. This will lead to a new approach to some of the problems of cosmological structure formation.
  11. 11.
    Bifurcations of Brans-Dicke cosmologies.
    The purpose here is to find the complete spectrum of bifurcations of Brans-Dicke cosmology for FRW models with a $\gamma$-law perfect fluid with or without a cosmological constant. The hyperbolic analysis has been done in various papers, but this project may also serve as an introduction to more complicated problems of a similar nature that arise in the context of effective string cosmology.
  12. 12.
    Bifurcation analysis of string cosmologies.
    One may think to revisit all five effective string theories and present qualitative analyses of flat FRW and Bianchi I cosmology in an effort to determine the nature of their different unfoldings. This analysis may delineate from another point of view the range dynamical behaviour of all five theories in these contexts and help to compare their predictions in concrete situations. Once one has this result, a study of the effects of other matter fields and spacetime geometries to better understand the nature of such predictions may thus be feasible.
  13. 13.
    Bifurcation theory of M-theoretic cosmology.
    This project studies 11-dimensional supergravity cosmologies compactified to 4-dimensional universes of various forms. For flat or open FRW universes a hyperbolic analysis has been done. A plan to apply in full the bifurcation sequence to a number of different 4-dimenisonal cosmologies in an M-theoretic context and compare the results to those of the string theory project may then be an interesting one.
  14. 14.
    Branes and bifurcations.
    We may reconsider the bulk-brane problem for 5-dimensional bulks with embedded de Sitter, anti-de Sitter or flat branes, for polytropic bulk fluids with or without a cosmological constant, taking into account the degenerate nature of the governing equations. Once one has constructed the possible bifurcation diagrams for this problem, properties of gravity localizing, bifurcating solutions on the brane which satisfy all energy conditions may be another project along these lines.
The last three projects are of a somewhat different nature than the previous ones. All projects discussed above may be considered to be in a sense a necessary step for this purpose.
  1. 1.
    The bifurcation sequence for the Schrödinger, Maxwell, and Dirac equations.
    Starting with calculations of the zero manifolds of the principal symbol in various cases like in vacuum and certain homogeneous media, one is hopefully led to a bifurcation treatment of the Schrödinger equation, the (linear) Maxwell or Dirac equations, and to further studies of topological invariants (like the Maslov index) that emerge in wave propagation in such contexts, and effects coming from the passage of rays through caustics.
  2. 2.
    Bifurcation theory with the full Einstein equations.
    It is a standing project to treat the full Einstein equations as a bifurcation problem and determine their (finite) codimension. The first step is to study the symbol of the Einstein equations and determine the zero manifold for a number of cases (vacuum and with matter). This is basically a problem of an algebraic-geometric nature.
  3. 3.
    Towards a (new?) fundamental theory.
    The versal unfolding equations modify the original set of gravity equations by adding new parameters and extra terms. Why is this modification physical rather than a contrived one? There is a ‘paradox’ involved here because the nonhyperbolic equilibria necessarily present in the original theory (say, general relativity), cannot really be treated properly unless one goes outside that theory and into one that is apparently qualitatively different than others and described by the versal unfolding construction.
    In other words, if one insists on studying any original theory per se, one must restrict their studies to only the ‘hyperbolic’ features of it. On the other hand, the versal unfolding of any equation (say the Friedmann equation, or that of the OS problem) exhausts in a sense everything that is to know about it, it represents a kind of a ‘theory of everything’ of that equation. Since all of the original equations belong to one and the same theory as special cases/problems of its full equations, the same must be true for their versal unfoldings. As discussed above, this seems to be in good correspondence with the principles set out in Ref. [93], p. 166. What is the nature of the structure that describes all these versal unfoldings [^6] We note that this interpretation of gravitational versal unfoldings is inequivalent to the one given here, say, for the versal unfolding of the simplest Friedmann equations, as an entropic effect and far-from equilibrium evolution. This is basically an interpretation within the framework of general relativity. On the contrary, interpreting the set of all versal unfoldings as comprising a (possibly) consistent theory is something completely different.?
@@ 4 Discussion
This work describes the path from the large body of known and well-understood results related to hyperbolic-like approaches to symmetric reductions in the form of differential dynamical systems of the Einstein equations in general relativity, to a consistent structurally stable, ‘versal’ framework, a new but yet largely physically unchartered structure that emerges when unfolding general relativistic reduced systems such as the five problems described in this paper. Our approach to this problem is based on fundamental mathematical works in the subject of bifurcation theory, starting with H. Poincaré’s Thesis in the year 1879, developed further by A. A. Andronov, R. Thom, V. I. Arnold, later by M. Golubitski, I. Stewart, F. Takens, R. I. Bogdanov, and more recently by F. Dumortier, H. Zoladek, and many other mathematicians. In this connection, we make the further following remarks.
@@
Consider the Einstein equations with a cosmological constant $\Lambda$, thought of as a parameter in the sense advanced in this paper. This is clearly an unfolding of the vacuum Einstein equations, but is it a versal unfolding of them [^7] An answer to this question could perhaps replace the question mark at the end of the vacuum equations in a famous photo!? We have shown that the Einstein equation with $\Lambda$ contains symmetric subsystems, e.g., the FL equations (with $\gamma=0$ in this case), for which the versal unfoldings have codimension greater than one, and three at most. The same is true for the evolution problem of event horizons, the crease flow also comes with codimension three when versally unfolded. There are further open questions related to these results, both mathematical as well as physical ones lying, however, outside the scope of the present paper and are presently at the level of pure speculation. A major problem in this direction is that we currently do not have any results on how the unfolded branches can be lifted by the full partial differential equation (we only know this in the zero-parameter case, and there only for the PDE-stability part of this question, not for the conjectured ‘centre manifold concentration’, which in fact may be false). For the physical direction, one may envisage the set of all versal unfoldings having a particular property, for example corresponding to homogeneous cosmologies. What is the nature of this set, and what is its physical significance? This set cannot correspond to a kind of modified gravity theory cosmology with additional action terms, because of the following observation. Consider the FRW problem discussed earlier in Section 3. In that problem, we start from the well-known results about the behaviour of the simplest FRW models and end up with their versal unfolding, namely, the equation $\Omega^{\prime}=\Omega(\Omega-1)+\nu$, where $\nu$ is the unfolding parameter. This is a perturbation of the Friedmann equations via adding lower-order terms (lower with respect to the first nonzero term in the Taylor expansion), and it is versal i.e., all terms of $O(3)$ or higher have no effect on the nature of the bifurcations. As we discussed, the solutions coming from the versal unfolding equation have entropy-related properties not met in the standard theory and emerge only as bifurcations. In this sense they represent not modifications of the standard theory but a certain completion of it at a deeper level (as we already discussed earlier), one that is very difficult to capture without bifurcation theory. This way gravitational bifurcation theory may considerably expand the present horizon of gravity research.
@@
Obviously one may consider unfolding a theory that contains higher-order curvature terms, or one that is of higher-dimensionality than four, for example in the framework of some modified gravity theory of effective string theory. Although this would lead to interesting projects in gravitational bifurcation theory, we note that there are already some answers to such a problem stemming from the results discussed in this paper. For example, suppose that one considers a Friedmann universe with a perfect fluid with or without a $\Lambda$ in a modified gravity or string theoretic context. If the linear part of the resulting equations turns out to be the same as the corresponding ones discussed here, then the answer will be identical to the ones we found here. This will be so due to the power of the versal unfolding construction. However, the inclusion of other matter fields, for instance a scalar field component, will drastically alter the conclusions and possibly increase the dimensionality of the dynamical systems and/or the codimension of the unfoldings (even in cases with identical symmetries as the ones here, for instance $\mathbb{Z}_{2}$). Indeed, developing the versal unfolding construction for a simple modified gravity theory in, say, a Friedmann context would be an interesting problem.
@@
Amongst further fundamental issues emerging in the framework of gravitational bifurcation theory and deserve a comment, we shall briefly discuss only three. Firstly, since the two appear intertwined in any bifurcation diagram, what is the relation between the unfolding parameter (‘motion’ in parameter space) and the time (evolution in phase space)? In particular, which of the two $\mu$ or $t$ is prior? Unconstrained motion in parameter space appears to be a cause for the appearance of a ‘time succession’ in phase space (metamorphoses). Consider evolution in a stratum without equilibria, e.g., the negative axis in Fig. 1 right, or the $\chi$-stratum in Fig. 4, and the motion of a parameter point around the origin in parameter space. Passing through a bifurcation set changes the number and nature of equilibria, and forces the phase point to move on a new orbit in the ‘next’ phase portrait. Such a smooth transition of the phase point between different phase portraits introduces a time progression into the problem associated with a global change (i.e., the appearance of one or more new equilibria). This state of affairs begs the question as to what causes motion in parameter space? We believe that an answer to this question might be related to the proven stability of the versal unfolding. In other words, the system collapses when the stability of the whole construction is disrupted, something which as we have shown cannot really happen in a versal unfolding.
Secondly, establishing a ‘general relativity theory landscape’ via gravitational bifurcation theory would not just be a formal curiosity, it provides a unifying framework that organizes the full zoo of known and yet-to-be-discovered solutions under one geometric roof (see Table 1). While Table 1 summarizes the deep parallels between string-theory and GR-bifurcation landscapes, the rigorous derivations of each entry are beyond the scope of this review and will be presented in forthcoming work. Here we highlight just a few key features of this correspondence. First, a versal-unfolding landscape endows GR with a moduli space of all symmetry-reduced solutions of Einstein’s equations, whose strata encode precisely how equilibrium branches appear, merge, or vanish as control parameters vary. This construction brings systematic control of gravitational instabilities: every boundary in parameter space marks a bifurcation at which a solution gains or loses stability, ensuring that no qualitative transition is overlooked. Second, the landscape approach bridges traditional PDE-level stability analyses and finite-dimensional ODE dynamics, extending linear results (e.g. around Kerr or FLRW) into a fully nonlinear classification of nearby behaviour and potentially uncovering new invariants or ‘charges’ associated with each stratum. Third, it forges powerful connections to algebraic-geometric and topological methods, such as discriminant varieties, resolution of singularities, and Hodge theory, providing tools to probe the global structure of solution spaces and phenomena such as crease flows on horizons. Finally, in direct analogy to string phenomenology, gravitational physicists can exploit this bifurcation landscape to guide numerical simulations and observational searches toward those regions of parameter space where novel phenomena (critical scaling laws, horizon echoes tied to global bifurcations, etc.) are most likely to emerge.
String-landscape concept GR-bifurcation-landscape analogue Comments
Moduli fields Control parameters (knobs) in a versal unfolding Codimension = number of knobs.
Vacua minima of $V(\phi)$ Equilibrium branches $x(\lambda)$ solving $f(x,\lambda)=0$ Stable when hyperbolic.
Scalar potential $V(\phi)$ Bifurcation function $f(x,\lambda)$ Zeros define equilibria.
Flux quantum numbers Codimension (number of degeneracy directions) Discrete knob choices.
Stability $\nabla^{2}V>0$ Hyperbolicity (invertible Jacobian) Structural-stability condition.
Tunneling Local bifurcations (branch creation/annihilation) Vacuum-decay analogue.
Discriminant loci Bifurcation surfaces in parameter space Boundaries of stability regions.
Metastability Persistence of branches under small changes Robustness measure.
Anthropic selection & dynamical selection Basin-of-attraction volumes Attractor preference.
Calabi-Yau moduli Versal-unfolding family Universal deformation space.
Flat (zero-frequency) directions Center manifold Slow-mode subspace.
Transition rates computed via instantons Normal-form analysis of bifurcations Semi-classical tunneling $\leftrightarrow$ normal form coefficients governing branch splitting.
Vacuum-tracking / phase diagram of vacua vs. moduli Bifurcation diagram Tracks branch splitting/merging.
Walls of marginal stability in moduli space Global bifurcations Large-scale separatrix reconnections.
Table 1: Analogy between key concepts in the string-theory landscape and their counterparts in the gravitational bifurcation landscape.